OpenStudy (anonymous):

Find the solution of this system of equations -3x-7y=-66 -10x-7y=-24

OpenStudy (anonymous):

First you probably just want to multiply either the top or bottom by -1

OpenStudy (igreen):

ILY?

OpenStudy (anonymous):

Then you can use elimination

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Lol @iGreen who are you talking to and do you still need my help @usitdown or can you take it from here?

OpenStudy (anonymous):

Still need help @ihateschool18

OpenStudy (igreen):

@ihateschool18 I'm talking to @usitdown he posted 'ILY @iGreen' to me..

OpenStudy (igreen):

Then he deleted it..

OpenStudy (anonymous):

Okay so I just multiplied the first equation by negative one so now we have \[\frac{ 3x+7y=66 }{ -10x-7y=-24 }\] so after elimination we are left with -7x=42. So x=-6.

OpenStudy (igreen):

Nice work @ihateschool18

OpenStudy (anonymous):

Now we plug x back into one of the original equation to get y.

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Is there a y value???

OpenStudy (igreen):

Yes, you plug it back into the equation.

OpenStudy (anonymous):

@AnswerMyQuestions

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

@AnswerMyQuestions !!!

OpenStudy (igreen):

\(-3x-7y=-66\) \(-3(-6)-7y=-66\) \(18-7y=-66\) \(-7y=-84\) \(y = 12\)

OpenStudy (igreen):

So your answer is (-6, 12) :)

OpenStudy (igreen):

@iGreen

OpenStudy (anonymous):

TY SO MUCH @iGreen . You are the smartest dude i know

OpenStudy (anonymous):

@iGreen I just bumped a question would you mind seeing if you could help

OpenStudy (anonymous):

Wow I feel loved @usitdown

OpenStudy (igreen):

LOL! The reason why he says that is because I helped him at many questions before..@ihateschool18

OpenStudy (igreen):

@ihateschool18 *

OpenStudy (igreen):

@IGREEN @iGreen @iGrEeN @igreen